OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = 5/2x + 2

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

the line \(y=\frac{5}{2}x+2\) has slope \(\frac{5}{2}\) the perpendicular line will therefore have slope \(-\frac{2}{5}\) the "negative reciprocal

OpenStudy (anonymous):

So the slope is -2/5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So would it be y=-2/5x-3

OpenStudy (anonymous):

??